OpenStudy (anonymous):

TRUE OR FALSE :The following is a geometric sequence. 1.5, 2.25, 3.375, 5.0625,…

OpenStudy (precal):

can you find a number that was multiplied or divided by all the way through? if so, then it is geometric

OpenStudy (anonymous):

See if consecutive terms have a common ratio.

OpenStudy (anonymous):

In other words,\[\large \frac{2.25}{1.5}\overset{?}{=}\frac{3.375}{2.25}\overset{?}{=}\frac{5.0625}{3.375}\overset{?}{=}\cdots\]If you don't have a calculator, try converting the decimals to fractions.